A lawn mower has a flat, rod-shaped steel blade that rotates about its center. The mass of the blade is 0.65 kg and its length is 0.55 m. What is the rotational energy of the blade at its operating angular speed of 3490 rpm ? If all of the rotational kinetic energy of the blade could be converted to gravitational potential energy, to what height would the blade rise?
Answer #1

A)

Rotational kE

= 0.5 I w^2

= 0.5* (1/12) m L^2 * w^2

= 0.5* ( 1/12)* 0.65* 0.55^2* (3490* 2* 3.14 / 60)^2

= 1093.19 J

=====

B)

Mgh = 1093.19

h = 1093.19 / ( 9.8* 0.65)

h = 171.615 m

======

Comment before rate in case any doubt, will reply for sure.. Goodluck